Circular No. 6466

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 N1 (BREWINGTON) Continuation to the ephemeris on IAUC 6435 (elements MPC 27690):

1996TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Sept. 4 14^{h}57.^{m}08 +50^{o}58.^{′}6 0.961 1.080 66.^{o}5 58.^{o}9  8.2
9 15 20.78 +53 09.2 0.967 1.127  69.5 56.9  8.4
14 15 47.22 +55 00.8 0.974 1.177  72.9 54.8  8.6
19 16 16.65 +56 30.1 0.981 1.230  76.6 52.6  8.9
24 16 48.98 +57 33.1 0.989 1.286  80.4 50.3  9.1
29 17 23.71 +58 05.6 0.999 1.343  84.4 47.9  9.3
Oct. 4 17 59.89 +58 04.0 1.012 1.403  88.4 45.5  9.5
9 18 36.28 +57 27.1 1.030 1.463  92.3 43.0  9.7
14 19 11.54 +56 16.3 1.054 1.525  96.0 40.6  9.9
19 19 44.63 +54 36.2 1.083 1.587  99.4 38.3 10.2
24 20 14.91 +52 33.4 1.120 1.649 102.3 36.1 10.4
29 20 42.17 +50 14.9 1.164 1.712 104.8 34.1 10.7
Nov. 3 21 06.50 +47 47.9 1.216 1.775 106.6 32.4 10.9

Total visual magnitude estimates: Aug. 13.87 UT, 7.5 (J. M. San Juan, Teruel, Spain, 10 × 70 binoculars); 16.86, 7.5 (M. V. Zanotta, Mt. Ponna, Italy, 7 × 42 binoculars); 19.89, 8.1 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); 22.89, 8.4 (A. Diepvens, Balen, Belgium, 0.15-m refractor); 25.83, 8.3 (K. Hornoch, Lelekovice, Czech Republic, 10 × 80 binoculars; moonlight).

1996 PW D. L. Rabinowitz, Jet Propulsion Laboratory, points out that his data

on IAUC 6452 gave reflectance colors, implying that the spectrum is indeed redder than solar (see also IAUC 6462). A refined reduction yields V = 17.42 ± 0.03 on Aug. 14.4 UT and 17.22 ± 0.07 on Aug. 15.4, with mean reflectance colors (from five observations) V − R = +0.18 ± 0.04, V − I = +0.34 ± 0.03, consistent with the spectra of S- or D-type objects.

NOVA CRUCIS 1996 Visual magnitude estimates extracted from VSNET, Kyoto, suggest

that this object (IAUC 6463) has faded rapidly: Aug. 28.702 UT, 10.6 (B. Monard, Pretoria, South Africa); 29.720, 10.8 (Monard); 30.429, 10.8 (P. Nelson, Ellinbank, N.S.W., Australia); 30.451, 11.0 (F. Farrell, Adelaide, S.A., Australia); 30.717, 11.3 (Monard).

1996 August 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden